An L-C circuit has an inductance of 0.350 HH and a capacitance of 0.290 nF . During the current oscillations, the maximum current in the inductor is 2.00 A .

Part A

What is the maximum energy EmaxEmaxE_max stored in the capacitor at any time during the current oscillations?

Express your answer in joules.

View Available Hint(s)

EmaxEmaxE_max =

nothing

JJ

Submit

Part B

How many times per second does the capacitor contain the amount of energy found in part A?

Express your answer in times per second.

View Available Hint(s)

nothing

s−1
Answer #1

A)

Energy, E = 0.5 LI^2

E = 0.5 x 0.35 x 2^2 = 0.7 J

B)

Angular frequency, w = 1/sqrt(LC) = 1/sqrt(0.35 x 0.29 x 10^-9) = 9.93 x 10^4 rad/sec

Time, T = 2 pi/w = 6.33 x 10^-5 sec

Since, 1 time period capacitor contains maximum energy 2 times.

So, in one second capacitor contains maximum energy

= 2/T = 3.16 x 10^4 times/sec

Comment in case any doubt please rate my answer....

​​​​​​